OpenStudy (anonymous):

3(x + 1) = −2(x − 1) + 6

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

3x + 3 = -2x +2 +6 3x + 2x = 2 + 6 - 3 5x = 5 x = 1

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome